Assuming all of the distributions are normal, find the Z-score for each of the cases:

a. µ = 17, σ = 5, x = 20, Z =?

b. µ = 365, σ = 14, x = 401, Z =?

c. µ = −11.94, σ = 0.73, x = −11, Z =?

d. (A-Grade) µ = 6.4, σ = −1.04, x = −5.2, Z =?
Answer #1

Solution:

Z = [x - µ ]/σ

a)

Z = [20 - 17]/5 = 0.60

b)

Z = [401 - 365]/14 = 2.57

c)

Z = [-11 - (−11.94)]/0.73 = 1.29

d)

Z = [ −5.2 - 6.4]/1.04 = -11.15